OpenStudy (anonymous):

based on "p^2-q^2=4" and "p-q=-1" which is bigger, "p+q" or "4"?

OpenStudy (anonymous):

\[p^2-q^2=(p-q)(p+q)=4\] and \[p-q=-1\implies p+q=-4\]

OpenStudy (anonymous):

Oh, I see how you did it, thanx!